### v1

# Influence of initial states on memory effects: A study of early-time superradiance

Abstract
The initial state of a quantum system can significantly influence its future dynamics, especially in non-Markovain quantum processes due to the environmental memory effects. Based on a previous work of ours, we propose a method to quantify the memory effects of a non-Markovian quantum process conditioned on a particular system initial state. We apply our method to study the early-time dynamics of a superradiance model where $N$ atoms (the system) interacting with a single-mode vacuum cavity (the environment) with several types of initial states. We find that the value of the memory effects in the early-time regime is half the environmental photon number for the (dephased) Dicke states. Besides, the memory effects, the environmental photon number and the degree of superradiance can be simultaneously enhanced by the coherence or entanglement of some initial states. In our study, the transitions from non-superradiant initial states to superradiant ones are always accompanied by the enhancement of memory effects, showing the importance of memory effects in superradiance.
I Introduction
The initial state of a quantum system can significantly influence its future dynamics [1, 2], especially in non-Markovain quantum processes due to the environmental memory effects. One trivial example is that if the system is initially in a steady state in a non-Markovian quantum process, it can hardly exhibit any non-Markovain features afterward, such as the revival of energy or information [3, 4]. More intriguing phenomena may emerge when a system consists of a collection of subsystems, such that the properties of its initial state, such as entanglement, may significantly influence the future dynamics. A well-known example is the concept of superradiance [5, 7, 6, 8, 9, 10, 11, 12] introduced by Dicke in 1954, where the emission intensity from an ensemble of atoms interacting with a common electromagnetic field can be enhanced compared with that from independent atoms. It is also well-known that the superradiance behaviors are highly relevant to the initial states of the system. In recent years, superradiance has received a large amount of attention due to its theoretical significance and potential applications [14, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Under certain approximations, such as a coarse grained time scale, the superradiance process could be regarded as Markovian [8, 9, 10, 11, 12], whereas, it is intrinsically non-Markovian. With the advances in theories and technologies, understanding the non-Markovian dynamics of superradiance becomes more demanding [26, 27, 28, 29]. In a recent work [29], the author shows that non-Markovian memory effects play an important role in superradiance beyond retardation, featuring the quadratic dynamics in the early-time (Zeno regime). In view of the significant influences of the system initial states on non-Markovain quantum processes, especially, a superradiance process, some interesting questions arise. For example, how to quantitatively evaluate the memory effects of a quantum process conditioned on a particular system initial state? How does the initial state of a system influence the memory effects? What is the role of the memory effects in a superradiance process, especially in its the early-time dynamics where the superradiance is created. Are there quantitative relations between the memory effects and the superradiance characteristics.
In this paper, we try to answer these questions by first proposing a method to quantify the memory effects in a quantum process conditioned on a particular system initial state. The method is based on a previous work [30] of ours. In that work, we quantify the memory effects (past-future dependence) of a quantum process through the inequality of completely positive dynamical maps $T(t_{2},t_{0})\neq T(t_{2},t_{1})T(t_{1},t_{0})$. By acting the inequality on a particular initial state $\rho_{S}(t_{0})$, the memory effects could be understood by the difference between two final states $\rho_{S}(t_{2})$ and $\rho_{S}^{\prime}(t_{2})$. The reason is that the evolutions corresponding to $T(t_{2},t_{0})$ and $T(t_{2},t_{1})$ have the same system states at $t_{1}$ but different histories before $t_{1}$. Based on the above interpretation, we suggest measuring the memory effects conditioned on the initial state $\rho_{S}(t_{0})$ by the maximal difference of $\rho_{S}(t_{2})$ and $\rho_{S}^{\prime}(t_{2})$ in a time interval of interest. Using this method, we calculate the memory effects (as well as the radiation characteristics) of a superradiance model in the early-time regime (Zeno regime). The model describes $N$ two-level atoms interacting with a single-mode vacuum cavity. Several types of initial states are studied, such as the (dephased) Dicke states and factorized identical states. We find that the value of memory effects for the (dephased) Dicke states is half the cavity photon number in the early-time regime. Besides, it is observed that the single-atom coherence in the factorized identical states and the coherence (entanglement) in the dephased Dicke states can enhance the memory effects, the cavity photon number and the degree of superradiance at the same time. In our studies, the transitions from non-superradiant initial states to superradiant ones are always accompanied by the enhancement of memory effects. The results demonstrate that the memory effects is vital for the superradiance phenomenon and the (change of) environment photon number is an important source of the memory effects, especially for the Dicke states.
The paper is organized as follows. In Sec.II, we propose a method to evaluate the initial-state-dependent memory effects in a quantum process. In Sec.III, we obtain the early-time solution of $N$ two-level atoms interacting with a vacuum cavity, after which we give the expressions for the memory effects, the cavity photon number and the degree of superradiance for an arbitrary initial state of the system. In Sec.IV, the influences of initial states on the memory effects as well as the superradiance characteristics are calculated and analyzed for several types of initial states. Conclusions and discussions are presented in Sec.V.
II Initial-state-dependent memory effects
In this section, we first review the measure of non-Markovianity in [30] that quantifies the memory effects (the past-future dependence) in a quantum process. Using its physical interpretations, we then propose a method to quantify the memory effects in a quantum process conditioned on a particular initial state of the system. The object of study in [30] is a quantum process of an open quantum system described by the total Hamiltonian
$H=H_{S}+H_{E}+H_{SE},$ (1)
an arbitrary system initial state $\rho_{S}(t_{I})$ and a fixed initial state of the environment $\rho_{E}(t_{I})$ (independent of the system). Here $t_{I}$ is the initial time of an evolution which is also arbitrary. The initial condition of an evolution is
$\rho_{SE}(t_{I})=\rho_{S}(t_{I})\otimes\rho_{E}(t_{I})$ (2)
which means that the system is initially uncorrelated with the environment before the evolution. In general, $\rho_{E}(t_{I})$ is governed by [31]
[…]
Typically, one deals with a time-independent environment Hamiltonian $H_{E}$ and assumes a steady state of $H_{E}$ as the environmental initial state (e.g., a thermal state). Then, the initial condition of an evolution is
$\rho_{SE}(t_{I})=\rho_{S}(t_{I})\otimes\rho_{E}$ (4)
for any $t_{I}$. Since $\rho_{S}(t_{I})$ and $t_{I}$ is arbitrary, the measure of non-Markovianity (memory effects) in [30] is determined by $H$ and $\rho_{E}(t_{I})$ and the time interval of interest.
In a non-Markovian quantum process, the meaning of a dynamical map given by $\rho_{S}(t_{2})=\varepsilon(t_{2},t_{1})\rho_{S}(t_{1})$ is ambiguous unless $\rho_{SE}(t_{1})$, or from another perspective, the initial time $t_{I}$ is specified ($t_{I}\leqslant t_{1}$). For clarity, we define $T(t_{2},t_{1})$ as a memoryless dynamical map that transfers $\rho_{S}(t_{1})$ to $\rho_{S}(t_{2})$ where $t_{1}$ is the initial time of the evolution ($t_{I}=t_{1}$) [30], i.e.,
$\rho_{S}(t_{2})$ $=$ $T(t_{2},t_{1})\rho_{S}(t_{1})$ (5)
$=$ $\text{Tr}_{E}[U(t_{2},t_{1})\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})U(t_{2},t_{1})^{{\dagger}}]$
[…]
$\rho_{S}(t_{2})$ $=$ $T(t_{2},t_{1})\rho_{S}(t_{1})$ (6)
$=$ $\text{Tr}_{E}[e^{-iH(t_{2}-t_{1})}\rho_{S}(t_{1})\otimes\rho_{E}e^{iH(t_{2}-t_{1})}]$
such that $T(t_{2},t_{1})=T(t_{2}-t_{1},0)$ [32]. The dynamical map $T$ is trace-preserving and completely positive and called a universal dynamical map (UDM) that is independent of the state it acts upon [33].
Let $t_{0}\leqslant t_{1}\leqslant t_{2}$, a (time-dependent) quantum process is Markovian if the dynamical maps (denoted by $\Lambda_{M}$) satisfy the divisibility condition
$\Lambda_{M}(t_{2},t_{0})=\Lambda_{M}(t_{2},t_{1})\Lambda_{M}(t_{1},t_{0})$ (7)
where each dynamical map is uniquely defined and a UDM. Remark that the dynamical map $\Lambda_{M}(t_{2},t_{1})$ is a UDM if and only if it is induced by
$\rho_{S}(t_{2})$ $=$ $\Lambda_{M}(t_{2},t_{1})\rho_{S}(t_{1})$ (8)
$=$ $\text{Tr}_{E}[U(t_{2},t_{1})\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})U(t_{2},t_{1})^{{\dagger}}]$
where $\rho_{E}(t_{1})$ is fixed and $\rho_{S}(t_{1})$ is arbitrary [33]. In a open quantum system, the condition $\rho_{SE}(t)=\rho_{S}(t)\otimes\rho_{E}(t)$ [$\rho_{E}(t)$ does not dependent on the system state] may not be satisfied exactly for $t>t_{I}$. Thus the dynamics of an exact open quantum system is typically not Markovian [33]. However, $\rho_{SE}(t)\approx\rho_{S}(t)\otimes\rho_{E}(t)$ can be a good approximation where the correlation between the system and the environment does not affect the system’s dynamics so much [33]. It is observed that the Markovian dynamical map $\Lambda_{M}(t_{2},t_{1})$ is unique and does not depend on the initial time of a evolution, e.g., $t_{I}=t_{1}$ or $t_{I}<t_{1}$. Therefore, $\Lambda_{M}(t_{2},t_{1})=T(t_{2},t_{1})$ according to the definition of the memoryless dynamical map $T$. Then, the Markovian divisibility condition Eq.(7) can also be expressed in terms of the memoryless dynamical map $T$ by
$T(t_{2},t_{0})=T(t_{2},t_{1})T(t_{1},t_{0})$ (9)
whose violation is a sign of non-Markovianity. Unlike the dynamical maps used in some non-Markovianity measures [34, 35], all the dynamical maps $T$ are completely positive. The violation of Eq.(9) is manifested by the inequality
$T(t_{2},t_{0})\neq T(t_{2},t_{1})T(t_{1},t_{0}).$ (10)
The physical meaning of Eq.(10) can be explained with Fig.1. Let the left-hand and the right-hand side of Eq.(10) act on a system initial state $\rho_{S}(t_{0})$. On the left-hand side of Eq.(10), $\rho_{S}(t_{0})$ is mapped to $\rho_{S}(t_{2})$ by $T(t_{2},t_{0})$ in evolution A. On the right-hand side, $\rho_{S}(t_{0})$ is firstly mapped to $\rho_{S}(t_{1})$ by $T(t_{1},t_{0})$ in evolution B. Then, as the initial state of evolution C, $\rho_{S}(t_{1})$ is mapped to $\rho^{\prime}_{S}(t_{2})$ by $T(t_{2},t_{1})$, which means that the initial condition of evolution C is $\rho_{SE}(t_{1})=\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})$ with fixed $\rho_{E}(t_{1})$ defined by Eq.(3) or a steady one $\rho_{E}$. At the moment $t_{1}$, evolution A and C have the same system state $\rho_{S}(t_{1})$ but different histories: evolution A has a history in time interval $[t_{0},t_{1}]$ that is encoded in $\rho_{SE}(t_{1})=U(t_{1},t_{0})\rho_{S}(t_{0})\otimes\rho_{E}(t_{0})U(t_{1},t_{0})^{{\dagger}}$; while evolution C (starting at $t_{1}$) have no history before $t_{1}$. Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ is an evidence that the future state (after $t_{1}$) of the system depends on its history (in $[t_{0},t_{1}]$) in this quantum process. This is the physical meaning of the memory effects in this paper and [30]. Otherwise, if the divisibility Eq.(9) holds, the process is Markovain and $\rho_{S}(t_{2})=\rho^{\prime}_{S}(t_{2})$ for any $\rho_{S}(t_{0})$.
On the other hand, the inequality could be understood by focusing on the change of the environment state between evolution B and C. At the end of evolution B, $\rho_{SE}(t_{1})=U(t_{1},t_{0})\rho_{S}(t_{0})\otimes\rho_{E}(t_{0})U(t_{1},t_{0})^{{\dagger}}$. Then, at the beginning of evolution C, the environment is initialized by $T(t_{2},t_{1})$ such that $\rho_{SE}(t_{1})\rightarrow\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})$ where $\rho_{E}(t_{1})$ is independent of the system. The information of the system’s history in $[t_{0},t_{1}]$ is erased by $T(t_{2},t_{1})$. In contrast, such an initialization never happens in evolution A. Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ signifies that the environment (as well as the correlations between the system and environment [33]) remembers the system’s history and this memory can influence the future of the system.
In [30], we define the maximal difference of $T(t_{2},t_{0})$ and $T(t_{2},t_{1})T(t_{1},t_{0})$ as the non-Markovianity where $t_{1}$ and $t_{2}$ are optimized. Based on the dynamical maps $T$, all the initial states are potentially considered, thus the measure of non-Markovianity in [30] does not depend on the initial state of the system. In this paper, our goal is to evaluate the influence of the system initial state on the memory effects in a quantum process. Using the physical interpretations discussed above, we define the value of memory effects conditioned on $\rho_{S}(t_{0})$ as the maximal trace distance between $\rho_{S}(t_{2})$ and $\rho^{\prime}_{S}(t_{2})$:
$N_{M}[\rho_{S}(t_{0})]=\max_{t_{1},t_{2}}\frac{1}{2}\|\rho_{S}(t_{2})-\rho^{\prime}_{S}(t_{2})\|.$ (11)
Here $||A||=\text{Tr}(\sqrt{A^{\dagger}A})$ is the trace norm of an operator $A$. Assuming a fixed $t_{0}$ for convenience, $N_{M}[\rho_{S}(t_{0})]$ could be calculated by optimizing $t_{1}$ and $t_{2}$ in a time interval $[t_{0},t]$ of interest where $t_{0}\leqslant t_{1}\leqslant t_{2}\leqslant t$. For example, the time interval might be $[t_{0},\infty]$ or $[t_{0},t_{0}+\tau]$ where $\tau$ is finite (as done later in this paper). Similar to the measure in [30], $0\leqslant N_{M}[\rho_{S}(t_{0})]\leqslant 1$ is satisfied due to the properties of the trace distance.
Using Eq.(11), the influence of initial states on the memory effects of an quantum process could be quantitatively evaluated. Note that $N_{M}[\rho_{S}(t_{0})]>0$ is a sufficient condition for the inequality Eq.(10), but not a necessary one. The theoretical calculation and experimental observation of Eq.(11) might be easier than those in [30] since the determination of the dynamical maps $T$ is not compulsory.
Figure 1: Physical interpretation of Eq.(10) as memory effects. At the moment $t_{1}$, the system states in evolution A and C are the same. Remind that evolution A has a history from $t_{0}$ to $t_{1}$ while evolution C does not have any history before $t_{1}$. Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ is an evidence that the future (after $t_{1}$) state of the system depends on its history (from $t_{0}$ to $t_{1}$) in a quantum process.
III Early-time superradiance
III.1 Theoretical model
We consider a fundamental model that describes $N$ two-level atoms (the system, denoted by “$S$”) interacting with a cavity (the environment, denoted by “$E$” ) initially in a vacuum state. The Hamiltonian is given by
$H\!$ $=$ $\!H_{S}+H_{E}+H_{I}$ (12)
$=$ $\!\omega_{A}\!\sum_{n=1}^{N}\!\sigma_{n}^{+}\sigma_{n}^{-}\!+\!\omega_{B}b^{{\dagger}}b\!+\!\sum_{n=1}^{N}g(\sigma_{n}^{+}b+\sigma_{n}^{-}b^{{\dagger}}).$
Here $H_{S}$ and $H_{E}$ represent $N$ noninteracting two-level atoms and a single-mode electromagnetic field in the cavity, respectively. $H_{I}=\sum_{n=1}^{N}g(\sigma_{n}^{+}b+\sigma_{n}^{-}b^{{\dagger}})$ describes the interactions between the atoms and the cavity with the rotating wave approximation, and the coupling strength $g$ is a real constant. The lowering and raising operators for the $n$th atom is defined as $\sigma_{n}^{-}=|g\rangle_{n}\langle e|_{n}$ and $\sigma_{n}^{+}=|e\rangle_{n}\langle g|_{n}$.
The initial condition of the model is assumed to be $\rho_{S\!E}(t_{0})=\rho_{S}(t_{0})\otimes|0\rangle\langle 0|$ where $\rho_{S}(t_{0})$ is arbitrary and $|0\rangle$ is the vacuum state of the cavity. For simplicity, we use $t_{0}=0$ in the reminder of this paper without loss of generality. The density matrix of the composite system $\rho_{S\!E}$ is described by the master equation,
$\dot{\rho}_{S\!E}$ $=$ $-i[H_{S}+H_{E}+H_{I},\rho_{S\!E}]$ (13)
$+\gamma(b\rho_{S\!E}b^{{\dagger}}-\frac{1}{2}b^{{\dagger}}b\rho_{S\!E}-\frac{1}{2}\rho_{S\!E}b^{{\dagger}}b)$
where $\gamma$ is the dissipation strength of the cavity. Assume the cavity is in resonance with the atoms ($\omega_{A}=\omega_{B}$), then the density matrix $\rho_{S\!E}$ in the interaction picture is described by
$\dot{\rho}_{S\!E}$ $=$ $-i[\sum_{n=1}^{N}g(\sigma_{n}^{+}b+\sigma_{n}^{-}b^{{\dagger}}),\rho_{S\!E}]$ (14)
$+\gamma(b\rho_{S\!E}b^{{\dagger}}-\frac{1}{2}b^{{\dagger}}b\rho_{S\!E}-\frac{1}{2}\rho_{S\!E}b^{{\dagger}}b).$
In this remainder of this paper, we focus on the early-time dynamics where $gt\ll 1$ due to the following reasons. First, non-Markovian characters are non-negligible on such a short time scale. Second, it stresses the influence of a particular initial state since the state hardly changes in this time interval. Furthermore, it is helpful for understanding the creation mechanism of the superradiance. In the early-time limit $gt\rightarrow 0$, one might ignore the influence of cavity dissipation $\gamma$, providing that $\gamma$ is not infinitely large. In this case, $\rho_{S\!E}$ evolves unitarily via
$\rho_{S\!E}(t)=U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t)$ (15)
where $U(t)=e^{-iH_{I}t}$ and $\rho_{E}=|0\rangle\langle 0|$. The influence of ignoring $\gamma$ on the early-time dynamics will be discussed at the end of this section.
III.2 Early-time solution
The reduced dynamics of the atoms could be determined by tracing out the degrees of the environment:
$\rho_{S}(t)$ $=$ $\text{Tr}_{E}[U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t)]$ (16)
$=$ $\sum_{k}\langle k|_{E}U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t)|k\rangle_{E}$
where $|k\rangle_{E}$ are a set of basis in $\mathcal{H}_{E}$. With the help of the Baker-Campbell-Hausdorff formula
$e^{\alpha A}Be^{-\alpha A}=B+\alpha[A,B]+\frac{\alpha^{2}}{2!}[A,[A,B]]+...,$ (17)
after the replacement $A=H_{I}$, $B=\rho_{S}(0)\otimes\rho_{E}$ and $\alpha=-it$, we have
$\rho_{S}(t)$$\approx$ $\sum_{k}\langle k|(\rho_{S}(0)\otimes\rho_{E}-it[H_{I},\rho_{S}(0)\otimes\rho_{E}]$ (18)
$-\frac{1}{2}t^{2}[H_{I},[H_{I},\rho_{S}(0)\otimes\rho_{E}]]|k\rangle$
where the terms with higher orders of $gt$ have been omitted due to the early-time limit ($gt\rightarrow 0$). Using the number state basis of the cavity $|k\rangle=|0\rangle,|1\rangle,...$ and the assumption $\rho_{E}=|0\rangle\langle 0|$, it is straightforward to derive the solution of the system evolution in the early-time limit $gt\rightarrow 0$,
$\rho_{S}(t)=\rho_{S}(0)+(gt)^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)].$ (19)
Here $\sigma^{-}=\sum_{n}\sigma_{n}^{-}$ is the collective lowering operator and $\mathcal{L}_{K}(\rho)=K\rho K^{+}-\frac{1}{2}K^{\dagger}K\rho-\frac{1}{2}\rho K^{\dagger}K$ is the Lindblad superoperator. $b^{{\dagger}}|0\rangle=|1\rangle$ and $\langle 0|b=\langle 1|$ have been used to obtain Eq.(19). Notice that the partial trace over the second term of Eq.(18) is zero, thus the change of $\rho_{S}(0)$ is quadratic in $t$ in the early-time limit.
III.3 Memory effects
With the above results, we now calculate the initial-state-dependent memory effects defined in Sec.II. We consider the dynamics in a short time interval $[0,t]$ where $t_{0}=0\leqslant t_{1}\leqslant t_{2}\leqslant t$. In “evolution A” (with a history from $t_{0}$ to $t_{1}$) as mentioned before, the initial state is $\rho_{S}(0)$ and the final state $\rho_{S}(t_{2})$ is given by
$\rho_{S}(t_{2})$ $=$ $\rho_{S}(0)+[g(\tau_{10}+\tau_{21})]^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)].$ (20)
Here $\tau_{10}=t_{1}-t_{0}$ and $\tau_{21}=t_{2}-t_{1}$ are used for convenience. The initial state of “evolution C” at $t_{1}$ is
$\rho_{S}(t_{1})$ $=$ $\rho_{S}(0)+(g\tau_{10})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)].$ (21)
In “evolution C” (without any history before $t_{1}$), the final state $\rho^{\prime}_{S}(t_{2})$ is given by
$\rho^{\prime}_{S}(t_{2})$ $=$ $\rho_{S}(t_{1})+(g\tau_{21})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(t_{1})].$ (22)
Substituting Eq.(21) into Eq.(22), we have
$\rho^{\prime}_{S}(t_{2})$ $=$ $\rho_{S}(0)+(g\tau_{10})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]$ (23)
$+(g\tau_{21})^{2}\mathcal{L}_{\sigma^{-}}\{\rho_{S}(0)+(g\tau_{10})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\}$
$=$ $\rho_{S}(0)+(g\tau_{10})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]$
$+(g\tau_{21})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]$
where the high-order term $g^{4}\tau_{10}^{2}\tau_{21}^{2}\mathcal{L}_{\sigma^{-}}\{\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\}$ has been omitted due to the early-time limit. The memory effects are manifested by the difference between Eq.(20) and Eq.(23):
$\rho_{S}(t_{2})\!-\!\rho^{\prime}_{S}(t_{2})\!$ $=$ $\!g^{2}[(\tau_{10}+\tau_{21})^{2}\!-\!\tau_{10}^{2}\!-\!\tau_{21}^{2}]\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]$ (24)
$=$ $\!2g^{2}\tau_{10}\tau_{21}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)].$
According to Eq.(11), the value of the memory effects for the initial state $\rho_{S}(0)$ is
$N_{M}[\rho_{S}(0)]$ $=$ $\max_{\tau_{10},\tau_{21}}\frac{1}{2}\|\rho_{S}(t_{2})-\rho^{\prime}_{S}(t_{2})\|$ (25)
$=$ $g^{2}\|\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\|\max_{\tau_{10},\tau_{21}}(\tau_{10}\tau_{21}).$
Since $g^{2}\|\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\|$ is a constant for a given $\rho_{S}(0)$, Eq.(25) can be calculated simply by maximizing $\tau_{10}\tau_{21}$ with the constraint $0\leqslant\tau_{10}+\tau_{21}=\tau_{20}\leqslant t$. Let $\tau_{10}+\tau_{21}=t^{\prime}$, then $\tau_{10}\tau_{21}=-(\tau_{10}-\frac{t^{\prime}}{2})^{2}+\frac{t^{\prime 2}}{4}\leqslant\frac{t^{\prime 2}}{4}$. Therefore, the maximum of $\tau_{10}\tau_{21}$ is $\frac{t^{\prime 2}}{4}$ when $\tau_{10}=\tau_{21}=\frac{t^{\prime}}{2}$. It is easy to see that the maximum of $\tau_{10}\tau_{21}$ in the time interval $[0,t]$ happens when $t^{\prime}=t$. Eventually, $N_{M}[\rho(t_{0})]$ in the early-time limit is
$N_{M}[\rho_{S}(0)]$ $=$ $g^{2}\|\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\|\max_{\tau_{10},\tau_{21}}(\tau_{10}\tau_{21})$ (26)
$=$ $g^{2}\|\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\|(\frac{t}{2})^{2}$
$=$ $\frac{1}{4}(gt)^{2}\|\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\|.$
Eq.(26) demonstrates that in a short time interval $[0,t]$, $N_{M}[\rho(0)]$ grows quadratically with $t$. To focus on the influence of initial states (rather than the time $t$ or the atom-cavity coupling $g$), it is convenient to discuss the normalized value of memory effects
$\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}=\frac{1}{4}\|\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\|.$ (27)
It represents the strength of memory effects in the early-time limit as a function of only $\rho_{S}(0)$. Notice that $\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}$ is not bounded as $N_{M}[\rho_{S}(0)]$. In principle, there is $0\leqslant\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}<\infty$.
III.4 Cavity photon number
In a non-Markovian proceess, the relation $\rho_{SE}(t)\approx\rho_{S}(t)\otimes\rho_{E}$ does not hold in general. Particularly, for the superradiance problem, the change of photo number in the environment (the cavity) might be an important source of the memory effects so that $\rho_{SE}(t)\neq\rho_{S}(t)\otimes\rho_{E}$. Thus it is desirable to know the cavity photon number (denoted by $N_{P}$ in this paper) in the early-time limit in order to understand the physics of the memory effects. The total excitation number of our model represented by $N_{ex}=b^{\dagger}b+\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}$ is conserved since $[N_{ex},H]=0$. Besides, there are no photon in the cavity initially. Therefore, the cavity photon number is equal to the loss of excitations of the atoms that represents the emission intensity of superradiance (in the early-time limit).
According to Eq.(19), the cavity photon number at $t$ for $\rho_{S}(0)$ is given by,
$N_{P}[\rho_{S}(0)]$ $=$ $\text{Tr}[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S}(0)]-\text{Tr}[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S}(t)]$ (28)
$=$ $\text{Tr}[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S}(0)]-\text{Tr}[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S}(0)]$
$-\text{Tr}\{\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}(gt)^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\}$
$=$ $(gt)^{2}\text{Tr}\{-\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\}.$
Using $[\sigma^{+}\sigma^{-},\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}]=0$ and $[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-},\sigma^{+}]=\sigma^{+}$, Eq.(28) can be simplified to a more concise form
$N_{P}[\rho_{S}(0)]=(gt)^{2}\text{Tr}[\sigma^{+}\sigma^{-}\rho_{S}(0)].$ (29)
It is observed that in the early-time limit, the cavity number increases quadratically with time. As mentioned in the last subsection, it is convenient to discuss the normalized cavity photon number
$\frac{N_{P}[\rho_{S}(0)]}{(gt)^{2}}=\text{Tr}[\sigma^{+}\sigma^{-}\rho_{S}(0)]$ (30)
such that it does not depend on the evolution time $t$ and atom-cavity coupling $g$. Eq.(30) represents the emission intensity in the early-time limit as a function of only the initial state. When each atom radiates independently, e.g., each atom radiates in its own cavity, the normalized cavity photon number for the $n$th atom is reduced to
$\frac{N_{P}^{(n)}[\rho_{S_{n}}(0)]}{(gt)^{2}}=\text{Tr}[\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S_{n}}(0)]$ (31)
where $\rho_{S_{n}}(0)$ is the reduced density matrix of the $n$th atom. Then the degree of superradiant might be measured by the ratio [7, 13]
$S[\rho_{S}(0)]$ $=$ $\frac{N_{P}[\rho_{S}(0)]/(gt)^{2}}{\sum_{n}N_{P}^{(n)}[\rho_{S}(0)]/(gt)^{2}}$ (32)
$=$ $\frac{\text{Tr}[\sigma^{+}\sigma^{-}\rho_{S}(0)]}{\sum_{n}\text{Tr}[\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S_{n}}(0)]}$
for nonzero denominator. If $S[\rho_{S}(0)]$ is greater (less) than 1, the state $\rho_{S}(0)$ is superradiant (subradiant).
III.5 Early-time regime
Mathematically, Eq.(19) holds true as $gt\rightarrow 0$. One may wonder the validity of the early-time solution Eq.(19) in a longer time interval and how the cavity dissipation deteriorates the validity. In this subsection, we discuss this problem by comparing the systems dynamics by Eq.(19) with that by numerically solving the full dynamics in $\mathcal{H}_{S}\otimes\mathcal{H}_{E}$ with Eq.(14) and tracing out the environment. A dynamics could be represented by the dynamical maps $T(t,0)$ that turns all the possible initial state at 00 to a final state at $t$. Furthermore, the difference of two dynamical maps can be evaluated through their Choi-Jamiółkowski matrices [36, 37]. Here we evaluate the error of Eq.(19) at instant $t$ by the trace distance of $\rho_{T(t,0)}^{quad}$ and $\rho_{T(t,0)}^{exact}$, i.e.,
$Error=\frac{1}{2}\|\rho_{T(t,0)}^{quad}-\rho_{T(t,0)}^{exact}\|$ (33)
where $\rho_{T(t,0)}^{quad}$ and $\rho_{T(t,0)}^{exact}$ are the Choi-Jamiółkowski matrices calculated by Eq.(19) and Eq.(14), respectively. The Choi-Jamiółkowski matrix is defined as $\rho_{T}(t,0)=T(t,0)\otimes\mathbb{I}(|\Psi\rangle\langle\Psi|)$ with $|\Psi\rangle=\frac{1}{\sqrt{2^{N}}}\sum_{i=1}^{2^{N}}|i\rangle|i\rangle$ a maximally entangled state of the $N$-atom system and an ancillary system of the same dimension.
We numerically calculated the error of Eq.(19) for $\gamma/g=0,1,10$ and $N=2,6$. The results are illustrated in Fig.2 in logarithmic scale. It is seen that the order of magnitude of error decreases linearly with that of $t$. Although the cavity dissipation and the increasing of atom number can increase the error, there exists a time interval where Eq.(19) is a good approximation as long as $\gamma$ and $N$ are finite. For example, in a time interval $[0,0.01/g]$, the error is less than $10^{-5}$ even when $\gamma=10g$ for both $N=2$ and $N=6$, implying that Eq.(19) is a good approximation in this scenario. The above discussion implies that the early-time limit $gt\rightarrow 0$ could be relaxed to an early-time regime, i.e., a time interval $[0,\tau]$ where the decay is quadratic. Here $\tau$ is linked to the Zeno time [38, 39, 40, 41].
Figure 2: The error of Eq.(19) as a function of $gt$ for (a) 2 atoms and (b) 6 atoms with different strengths of cavity dissipation. The error represents the difference of the dynamical maps $T(t,0)$ corresponding to Eq.(19) (early-time assumption) and Eq.(14) (exact). The error vanishes as $gt\rightarrow 0$ and increases with the cavity dissipation $\gamma$ or the atom number $N$.
IV Influence of initial states
In this section, the results Eq.(27), Eq.(30) and Eq.(32) are applied to several types of $N$-atom initial states. Our first aim is to study the influence of initial states on the memory effects. Another aim is to reveal the role of memory effect in creating superradiance. Conversely, it also helps to understand the physical source of memory effects in our model.
IV.1 Dicke states
The Dicke states, written as $|JM\rangle$, are extensively studied in the field of superradiance. It is defined as the common eigenstate of the pseudospin operators $D^{2}=\frac{1}{2}(D^{+}D^{-}+D^{-}D^{+})+D_{z}^{2}$ and $D_{z}$ with $J=N/2$. The eigenvalues are given by
$D^{2}|JM\rangle$ $=$ $J(J+1)|JM\rangle$ (34)
$D_{z}|JM\rangle$ $=$ $M|JM\rangle\qquad(M=-J,...,J)$ (35)
where the operators are defined by $D^{+}=\sum_{n}\sigma_{n}^{+}$, $D^{-}=\sum_{n}\sigma_{n}^{-}$ and $D_{z}=\frac{1}{2}\sum_{n}(|e\rangle_{n}\langle e|_{n}-|g\rangle_{n}\langle g|_{n})$. A Dicke state is symmetrical and invariant by atom permutation with $N_{e}=J+M$ excited atoms and $N_{g}=J-M$ ground-state atoms. It can be constructed by the following formula [8]:
$|JM\rangle=\sqrt{\frac{(J+M)!}{N!(J-M)!}}(\sum_{n}\sigma_{n}^{-})^{J-M}|e,e,...,e\rangle.$ (36)
For example, for 3 two-level atoms, there is $|3/2,-1/2\rangle=\frac{1}{\sqrt{3}}(|egg\rangle+|geg\rangle+|gge\rangle)$.
A well-known result of the free-space spontaneous emission rate for two-level atoms is $W_{N}=\Gamma\langle\sigma^{+}\sigma^{-}\rangle$ where $\Gamma$ is the natural linewidth of a single atom. For the Dicke states, there is
$W_{N}$ $=$ $\Gamma\langle JM|\sigma^{+}\sigma^{-}|JM\rangle$ (37)
$=$ $\Gamma(J+M)(J-M+1)$
When $M=J$ (fully excited), $W_{N}=N\Gamma$. Then the emission rate is proportional to $N$, or say, the $N$-atom emission rate is equal to the summation of those from independent atoms. Thus there is no superradiance. When $M=0$ (half excited), $W_{N}=\Gamma\frac{N}{2}(\frac{N}{2}+1)$, the emission rate increase with $N^{2}$, or say, the $N$-atom emission rate is greater than the summation of those from independent atoms ($\frac{N}{2}\Gamma$). The superradiance with the most intensive emission happens in this case. Note that the results holds under the Born-Markov approximation and on a coarse-grained time scale [8, 9, 10, 11, 12].
[…]
In the non-Markovian early-time regime, we calculated the normalized value of memory effects and cavity photon number with Eq.(27) and Eq.(30) for the Dicke states of different $N$. The results are shown in Fig.3 where $N=1,2,...,15$ ($J=N/2$) and $M=-J,-J+1,...,J$. It is observed that when the number of atoms $N$ is fixed, the strongest memory effects as well as photon emission happens when $M$ is 0 or next to 0, which is in agreement with Eq.(37). Interestingly, we find that the value of memory effects for the Dicke states can be fully determined by the cavity photon number by
$N_{M}(|JM\rangle)=\frac{1}{2}N_{P}(|JM\rangle)$ (38)
in the early-time regime. The relation is proved in the following. Using the property $\sigma^{\pm}|JM\rangle=\sqrt{J(J+1)-M(M\pm 1)}|J(M\pm 1)\rangle$, it is straightforward to see that
$\frac{N_{M}(|JM\rangle)}{(gt)^{2}}\!$ $=$ $\!\frac{1}{4}\|\mathcal{L}_{\sigma^{-}}(|JM\rangle)\|$ (39)
$=$ $\!\frac{1}{4}\|\sigma^{-}|JM\rangle\langle JM|\sigma^{+}\!-\!\frac{1}{2}\sigma^{+}\sigma^{-}|JM\rangle\langle JM|$
$\!-\frac{1}{2}|JM\rangle\langle JM|\sigma^{+}\sigma^{-}\|$
$=$ $\!\frac{1}{4}(J\!+\!M)(J\!-\!M\!+\!1)\big{\|}|JM^{\prime}\rangle\langle JM^{\prime}|$
$-|JM\rangle\langle JM|\big{\|}$
where $M^{\prime}=M-1$. Remind that the trace norm could be calculated by $\|A\|=\text{Tr}\sqrt{A^{\dagger}A}=\sum_{m}|\lambda_{m}|$ where $\lambda_{m}$ is the eigenvalue of operator $A$. Meanwhile, the eigenvalues of $|JM^{\prime}\rangle\langle JM^{\prime}|-|JM\rangle\langle JM|$ is $1$ and $-1$ whose absolute values sum up to 2. Therefore,
$\frac{N_{M}(|JM\rangle)}{(gt)^{2}}=\frac{1}{2}(J+M)(J-M+1).$ (40)
Considering that the normalized cavity photon number for Dicke states is
$\frac{N_{P}(|JM\rangle)}{(gt)^{2}}$ $=$ $\text{Tr}(\sigma^{+}\sigma^{-}|JM\rangle\langle JM|)$ (41)
$=$ $\text{Tr}[(J+M)(J-M+1)|JM\rangle\langle JM|]$
$=$ $(J+M)(J-M+1),$
Eq.(38) is proved. The relation Eq.(38) demonstrates that the (change of) the cavity photon number is a fundamental reason of the memory effects in the early-time regime for the Dicke states.
We now discuss the degree of superadiance for the Dicke states in the early-time regime. The normalized cavity photon number for one independent atom in its excited state is given by $N_{P}(|e\rangle)/(gt)^{2}=N_{P}(|\frac{1}{2},\frac{1}{2}\rangle)/(gt)^{2}=1$ by Eq.(31). Using the results of Eq.(41), the degree of superradiance for the Dicke states could be calculated by
$S(|JM\rangle)$ $=$ $\frac{N_{P}(|JM\rangle)/(gt)^{2}}{\sum_{n=1}^{J+M}N_{P}(|e\rangle)/(gt)^{2}}$ (42)
$=$ $\frac{(J+M)(J-M+1)}{J+M}$
$=$ $J-M+1$
for $J+M>0$, where the denominator represents the contribution of the emission from the $J+M$ independently excited atoms in $|JM\rangle$. Note that the physical meaning of the denominator in Eq.(42) is different from that in Eq.(32), but their values are the same.
In Fig.3(a) or (b), the left row represents $|J,-J\rangle$ (no atom excited), thus $N_{M}$ and $N_{P}$ are both zero which is a trivial case. Another group of Dicke states without superradiance are $|J,J\rangle$ (fully excited) corresponding to the right row in Fig.3(a) or (b). Using Eq.(42), it is easy to check that $S(|J,J\rangle)=1$. Except for the two rows, other Dicke states in Fig.3 are superradiant which could be examined by $S(|JM\rangle)>1$ with Eq.(42). For example, for the Dicke states $|J,-J+1\rangle$ (one atom excited) represented by the second row from the left, there is $S(|J,-J+1\rangle)=N$. The single-photon superradiance is called “the greatest radiation anomaly” by Dicke. Meanwhile, the normalized value of memory effects for $|J,-J+1\rangle$ is $\frac{N}{2}$, which is also proportional to $N$. For the states in the second row from the right, there is $S(|J,J-1\rangle)=2$ for $N>1$. The degree of superradiance is a constant showing a weaker degree of superradiance. For Dicke states, although the normalized value of memory effects is not proportional to the degree of superradiance in general, it is observed that the transition from non-supperradiant states to superradiant states (e.g., $|J,J\rangle\rightarrow|J,J-1\rangle$) is always accompanied by the increase of memory effects. Thus the memory effects are important for the creation of superradiance.
IV.2 Factorized identical states
In this subsection, we investigate $N$-atom initial states in a factorized from
$\rho_{S}^{fact}(0)$ $=$ $\rho_{S_{1}}\otimes\rho_{S_{2}}\otimes...\otimes\rho_{S_{N}}$ (43)
with identical single-atom states
$\rho_{S_{1}}=\rho_{S_{2}}=...=\rho_{S_{N}}=\left(\begin{array}[]{cc}\rho_{ee}&\rho_{eg}\\ \rho_{eg}^{*}&\rho_{gg}\\ \end{array}\right).$ (46)
Such states are fully characterized by $\rho_{ee}$ and $\rho_{eg}$ and may also be superradiant. For example, states of this form are experimentally studied in [13] to realize the single-atom superradiance where the a serious of atoms enter a cavity one by one. In [21], we show that such a model can be used to achieve the Heisenberg Limit in a quantum metrology scheme for measuring the atom-cavity coupling. The superradiance of two artificial atoms with initial state $\frac{1}{2}(|e\rangle+|g\rangle)\otimes(|e\rangle+|g\rangle)$ are experimentally observed in [15]. In the above works, the single-atom coherence ($|\rho_{eg}|$) are crucial for creating superradiance. Thus it is desirable to know the influence of such states on the memory effects in superradiance, especially the role of the single-atom coherence.
In this subsection, the normalized value of memory effects and the normalized cavity photon number for states described by Eq.(43) are calculated using Eq.(27) and Eq.(30). When Eq.(46) is a pure state, i.e., $\rho_{Sn}=|\phi\rangle\langle\phi|$, $\rho_{S}^{fact}(0)$ can be decomposed into the superpositions of different Dicke states [13]. Using the analytical result of $\langle\phi|\sigma^{+}\sigma^{-}|\phi\rangle$ in [13], we obtain the normalized cavity photon number for $\rho_{S}^{fact}(0)$:
$\frac{N_{P}[\rho_{S}^{fact}(0)]}{(gt)^{2}}=N(N-1)|\rho_{eg}|^{2}+N\rho_{ee}.$ (47)
According to Eq.(32), the degree of superradiance is given by
$S[\rho_{S}^{fact}(0)]$ $=$ $\frac{N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}}{\sum_{n=1}^{N}N_{P}^{(n)}(\rho_{S_{n}})/(gt)^{2}}$
$=$ $\frac{N(N-1)|\rho_{eg}|^{2}+N\rho_{ee}}{N\rho_{ee}}$
$=$ $1+(N-1)|\rho_{eg}|^{2}/\rho_{ee}$
for $\rho_{ee}>0$. The result demonstrates that the state $\rho_{S}^{fact}(0)$ is superradiant when $|\rho_{eg}|$ is nonzero and $N>1$. Besides, the normalized cavity photon number as well as the degree of superradiance grow quadratically with the single-atom coherence $|\rho_{eg}|$ for $N>1$.
Figure 4: Normalized memory effects $N_{M}$ (left column) and normalized cavity photon number $N_{p}$ (right column) for $\rho_{S}^{fact}(0)$ as a function of $\rho_{ee}$ and $|\rho_{eg}|$ in the early-time regime. The atom numbers are $N=1$, $2$ and $11$ from top to bottom. For any $\rho_{ee}$ and $N$ ($N>1$), $N_{M}$ ($N_{p}$) can always be enhanced nonlinearly by the single-atom coherence $|\rho_{eg}|$.
In contrast to Eq.(40), the analytical solution of Eq.(27) for the factorized initial states is complicated in general. Therefore, Eq.(27) is solved numerically in this section. The results of the normalized $N_{M}[\rho_{S}^{fact}(0)]$ and $N_{P}[\rho_{S}^{fact}(0)]$ are depicted in Fig.4 as a function of $\rho_{ee}$ and $|\rho_{eg}|$ with $N=1$, 2 and $11$. The condition $|\rho_{eg}|^{2}\leqslant(1-\rho_{ee})\rho_{ee}$ are satisfied in Fig.4 to guarantee the positivity of $\rho_{S_{n}}$. The single-atom coherence $\rho_{eg}$ are chosen to be real in the simulations for simplicity. Numerical results shows that the phase of $\rho_{eg}$ does not affect the value of memory effects. Unlike the case for the Dicke states, $N_{M}[\rho_{S}^{fact}(0)]/(gt)^{2}$ is not proportional to $N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}$ in general. However, our simulation results show that the normalized value of memory effects can always be enhanced by $|\rho_{eg}|$ for any $N$. Similar situations happen for $N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}$ and $S[\rho_{S}^{fact}(0)]$ when $N>1$ as seen in Eq.(47) and (IV.2). Besides, when the atom number $N$ increases, $N_{M}[\rho_{S}^{fact}(0)]/(gt)^{2}$ grows more and more fast with $|\rho_{eg}|$, and so do $N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}$ and $S[\rho_{S}^{fact}(0)]$ as seen in Eq.(47) and (IV.2). Thus we conclude that the creation and enhancement of superradiance by $|\rho_{eg}|$ is accompanied by the increase of memory effects for the factorized identical initial states. In Fig.4(c) and (e), it is found that the states around $\rho_{eg}=0$ and $\rho_{ee}=0.5$ tend to take lower values of memory effects compared with their neighboring states, especially when $N$ is large. This might be intuitively interpreted by the fact that $\rho_{eg}=0$ and $\rho_{ee}=0.5$ leads to a maximally mixed state of the $N$-atom system. Besides, Fig.4 shows that the (change of) the cavity photon number is not the only factor to determine the value of the memory effects. Other factors might include the (change of) coherence of the cavity state between the basis $|n\rangle$.
IV.3 Other initial states
Entanglement and coherence of the initial states may play important roles in superradiance [42, 22, 23, 24, 25], especially for the Dicke states. Next, we examine the influence of the coherence (or the entanglement among the atoms) in the dephased Dicke states $\rho^{deph}_{S}(0)$ on the memory effects as well as the superradiance. The initial state we consider is given by
$\rho^{deph}_{S}(0)=\lambda|JM\rangle\langle JM|+(1-\lambda)\mathcal{D}(|JM\rangle\langle JM|)$ (48)
where $0\leqslant\lambda\leqslant 1$. Here $\mathcal{D}$ turns $|JM\rangle\langle JM|$ into a fully dephased state with vanishing off-diagonal elements in the basis $|e(g),e(g),...,e(g)\rangle$. Meanwhile, $\mathcal{D}(|JM\rangle\langle JM|)$ is a mixed state of separable states whose entanglement is zero. Thus $\lambda$ reflects the strength of entanglement or the value of coherence of $\rho^{deph}_{S}(0)$. With the help of numerical solutions, we find by mathematical induction that $\frac{N_{M}[\mathcal{D}(|JM\rangle\langle JM|)]}{(gt)^{2}}=\frac{1}{2}(J+M)$ and
$\frac{N_{M}[\rho^{deph}_{S}(0)]}{(gt)^{2}}$ $=$ $\lambda\frac{N_{M}(|JM\rangle\langle JM|)}{(gt)^{2}}$ (49)
$+(1-\lambda)\frac{N_{M}[\mathcal{D}(|JM\rangle\langle JM|)]}{(gt)^{2}}$
$=$ $\frac{1}{2}(J+M)[(J-M)\lambda+1].$
It is seen that the normalized value of memory effects increases linearly with $\lambda$ when $M\neq J,-J$. The normalized cavity photon number can be calculated by Eq.(30). We obtain
$\frac{N_{P}[\rho^{deph}_{S}(0)]}{(gt)^{2}}=$ $=$ $\lambda\frac{N_{P}(|JM\rangle)}{(gt)^{2}}$ (50)
$+(1-\lambda)\frac{N_{P}[\mathcal{D}(|JM\rangle\langle JM|)]}{(gt)^{2}}$
$=$ $(J+M)[(J-M)\lambda+1].$
Similarly, $N_{M}[\rho^{deph}_{S}(0)]=\frac{1}{2}N_{P}[\rho^{deph}_{S}(0)]$ is satisfied for the dephased Dicke states in the early-time regime. The degree of superradiance for the dephased Dicke states is given by
$S[\rho^{deph}_{S}(0)]$ $=$ $\frac{N_{P}[\rho^{deph}_{S}(0)]/(gt)^{2}}{\sum_{n=1}^{J+M}N_{P}(|e\rangle)/(gt)^{2}}$ (51)
$=$ $\frac{(J+M)[(J-M)\lambda+1]}{J+M}$
$=$ $(J-M)\lambda+1$
for $J+M>0$. It is observed that the entanglement or coherence (represented by $\lambda$) in the depahsed Dicke state is necessary for superradiance. Besides, the degree of superradiance increases linearly with $\lambda$ for superradiant Dicke states. Similarly, the creation and increase of superradiance induced by $\lambda$ is accompanied by the increase of memory effects.
At the end of this subsection, we calculate the value of the normalized value of memory effects of a probabilistic mixture of Dicke states $\rho_{S}^{mix}(0)=\sum_{M}p_{M}|JM\rangle\langle JM|$ where $J=N/2$, $M=-J,-J+1,...,J$ and $\sum_{M}p_{M}=1$. Using similar treatment as in Eq.(39), we find that
$\frac{N_{M}[\rho_{S}^{mix}(0)]}{(gt)^{2}}\!$ $=$ $\!\frac{1}{4}[\!\sum_{M=-J}^{J-1}\!|p_{M+1}f(M+1)-p_{M}f(M)|$ (52)
$+p_{J}f(J)]$
where $f(X)=(J+X)(J-X+1)$. The result reduces to Eq.(40) when $\rho_{S}^{mix}(0)$ is a pure state $|JM\rangle\langle JM|$.
V Conclusions and Discussions
In this paper, we propose a method to evaluate the memory effects in a quantum process conditioned on a particular system initial state. The method is based on the physical interpretations of the inequality $T(t_{2},t_{0})\neq T(t_{2},t_{1})T(t_{1},t_{0})$ [30]. Some features of the non-Markovianity measure in [30] are inherited. For example, nonzero memory effects can be characterized even in regimes where monotonic behaviors do not occur, or have not occurred yet. Besides, $0\leqslant N_{M}[\rho_{S}(t_{0})]\leqslant 1$ is satisfied without additional normalizations. This allows us to compare the memory effects of quantum systems with different dimensions (e.g., different numbers of atoms in our model). Using our method, we calculate the influence of several types of initial states on the early-time memory effects. For the (dephased, mixed) Dicke states, we obtained simple analytic solutions for the normalized memory effects in terms of $J$ and $M$, or say, $N_{e}$ and $N_{g}$. The environmental photon number as well as the degree of superradiance are also calculated. The aims of comparing the memory effects with the superradiance characters is to reveal the role of memory effects in generating superradiance, and conversely, understand the physical sources of the memory effects.
The main conclusions are as follows: For the (dephased) Dicke states, the value of memory effects is half the cavity photon number in the early-time regime. For the factorized identical states, the single-atom coherence is necessary for superradiance. Such coherence can always enhance the memory effects for $N\geqslant 1$, meanwhile, it increases the cavity photon number and the degree of superradiance for $N\geqslant 2$. For the dephased Dicke states, the coherence (entanglement of atoms) represented by $\lambda$ is also necessary for superrdiance. Meanwhile, the memory effects, the cavity photon number and the degree of superradiance increase linearly with $\lambda$ for superradiant states. In the above cases, the transitions from non-superradiant states to superradiant ones are always accompanied by the enhancement of memory effects, implying the memory effects are vital for creating superradiance. On the other hand, we find that the (change of) environmental photon number is an important (but may not be the only) source of the memory effects in the superradiance process, especially for the Dicke states.
VI ACKNOWLEDGMENTS
The authors thank X. L. Huang, X. L. Zhao, J. Cheng and B. Cui for helpful discussions. The work is supported by the National Natural Science Foundation of China under Grant No. 11705026, the Fundamental Research Funds for the Central Universities in China under Grant No. 3132020178 and the Natural Science Foundation of Liaoning Province under Grant No. 2023-MS-333.